# On local Galois deformation rings: generalised reductive groups

### v1

@@ 13.4. Consequences
Let $x$ be either a closed point of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$ or a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}}$. Then the following hold:
@@
    $R^{\square}_{G,\rho_{x}}$ is a flat $\Lambda$-algebra of relative dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$ and is complete intersection;
@@
    if $\mathrm{char}(\kappa(x))=p$ then $R^{\square}_{G,\rho_{x}}/\varpi$ is complete intersection of dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$.
@@
The proof is essentially the same as the proof of [8, Corollary 3.38], so we only give a sketch. Corollary 3.8 and Lemma 3.9 applied with $A=\Lambda$ and $B=R^{\square}_{G,\rho_{x}}$ imply that to prove part (1) it is enough to show that
@@
If $\mathrm{char}(\kappa(x))=0$ then $\Lambda=\kappa(x)$. If $\mathrm{char}(\kappa(x))=p$ then $\Lambda$ is a DVR with uniformizer $\varpi$ and the fibre ring in (80) is $R^{\square}_{G,\rho_{x}}/\varpi$. Thus, if we can prove (80) then Lemma 3.9 applied with $A=\kappa(x)$ and $B=R^{\square}_{\rho_{x}}/\varpi$ will imply part (2).
@@
The proof is essentially the same as the proof of [8, Corollary 3.40] so we only give a sketch. Lemma 5.17 relates $R^{\square}_{G,\rho_{x}}$ and the local ring $\mathcal{O}_{X,x}$, where $X:=X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $x$ is a closed point of $X$. Using Corollary 13.27 we deduce that $\mathcal{O}_{X,x}$ are complete intersection and $\mathcal{O}$-flat for all closed points $x$. Thus $X$ is flat over $\operatorname{Spec}\mathcal{O}$ and locally complete intersection. The last property implies that $X$ is equidimensional. The same argument works for the special fibre. ∎
@@
We will prove (1); the proof of (2) is identical. Let $\eta$ be a generic point of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Let $\bar{\eta}$ be a geometric point above $\eta$. It follows from [52, Theorem 3 (ii)] that there is a geometric point $\xi:\operatorname{Spec}\kappa(\bar{\eta})\rightarrow X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ above $\bar{\eta}$. Let $X$ be an irreducible component of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ containing $\xi$ and let $X^{\mathrm{git}}$ be the image of $X$ in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Then $X^{\mathrm{git}}$ contains $\eta$. It follows from [8, Lemma 2.1] that $X$ is $G^{0}$-invariant. Since $X$ is closed in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, [52, Theorem 3 (iii)] implies that $X^{\mathrm{git}}$ is closed in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. If $X=\operatorname{Spec}A$ then $X^{\mathrm{git}}=\operatorname{Spec}A^{G^{0}}$, and since $X$ is irreducible, we deduce that $X^{\mathrm{git}}$ is irreducible. We conclude that $X^{\mathrm{git}}$ is the irreducible component of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ containing $\eta$.
@@
Corollaries 13.31 and 13.32 are the analogues of the main results of [9]; compare [9, Theorem 5.5.1]. Even if $G=\operatorname{GL}_{d}$ the proof given here is new: although we use many techniques introduced by Böckle and Juschka in [9], we never appeal to [9, Section 5].
@@
Using Lemma 3.10 it is enough to prove the statement after replacing $\kappa$ by a finite extension. Using Proposition 6.6 we may assume that there exist $g\in G^{0}(\kappa)$ such that $\rho^{\prime}(\gamma):=g\rho(\gamma)g^{-1}\in G(\mathcal{O}_{\kappa})$ for all $\gamma\in\Gamma_{F}$. Since $G$ is smooth over $\mathcal{O}$ we may pick $h\in G(\Lambda)$ which maps to $g$ in $G(\kappa)$. Conjugation by $h$ induces an isomorphism of $\Lambda$-algebras between $R^{\square}_{G,\rho}$ and $R^{\square}_{G,\overline{\rho}^{\prime}}$, so we may assume that $\rho$ takes values in $G(\mathcal{O}_{\kappa})$. Let $\overline{\rho}$ be the reduction of $\overline{\rho}$ modulo the uniformizer of $\mathcal{O}_{\kappa}$. Let $\mathcal{O}$ be the Witt ring of residue field of $\mathcal{O}_{\kappa}$. Since $\rho$ is a deformation of $\overline{\rho}$ to $\mathcal{O}_{\kappa}$ we obtain a map of local $\mathcal{O}$-algebras $R^{\square}_{G,\overline{\rho}}\rightarrow\mathcal{O}_{\kappa}$. Lemma 3.11 says that $R^{\square}_{G,\overline{\rho}}$ is naturally isomorphic to the completion of $\Lambda\otimes_{\mathcal{O}}R^{\square}_{G,\overline{\rho}}$ with respect to the kernel $\mathfrak{q}$ of the natural map to $\kappa(x)$. Corollary 13.27 together with [8, Lemmas 3.36, 3.37] allow us to bound the dimension of the completion by $\dim\Lambda+\dim G_{k}([F:\mathbb{Q}_{p}]+1)$, and if $\mathrm{char}(\kappa(x))=p$ then the dimension of the special fibre of the completion by $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$. We obtain $\dim\kappa\otimes_{\Lambda}R^{\square}_{G,\rho}\leq\dim G_{k}([F:\mathbb{Q}_{p}]+1)$ and the proof of Corollary 13.27 goes through. ∎
@@
Every continuous representation $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ can be lifted to characteristic zero.
@@
The proof is essentially the same as the proof of [8, Proposition 3.55]. The map $X^{\square}_{G,\overline{\rho}}\rightarrow X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is flat, since it is a localisation followed by completion. Let $V$ be a preimage of $V^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ in $X^{\square}_{G,\overline{\rho}}$. Then $V$ is contained in $X^{\square}_{G,\overline{\rho}}\setminus\{\ast\}$ and Lemma 3.54 in [8] together with Corollary 13.30 implies that $V$ is Zariski dense in $X^{\square}_{G,\overline{\rho}}$. The set of closed points in $V$ is equal to $\Sigma^{\operatorname{n-spcl}}$. Since $V$ is open $X^{\square}_{G,\overline{\rho}}\setminus\{\ast\}$, it is Jacobson, and hence $\Sigma^{\operatorname{n-spcl}}$ is dense in $V$ and also in $X^{\square}_{G,\overline{\rho}}$. The same argument works for the special fibre. ∎
@@ 14. Non-special locus
Let $H$ be a closed generalised reductive subgroup of $G$ such that $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ is non-empty. Then $H^{0}$ acts on $G^{0}\times X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ via $h\cdot(g,x)=(gh^{-1},h\cdot x)$. We let $G^{0}\times^{H^{0}}X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ be the GIT-quotient for this action. Since the action is free the fibre at a geometric point $x$ is isomorphic to $H^{0}_{\kappa(x)}$. This implies that
@@
The following Lemma formalises an argument made in [8, Proposition 3.8].
@@
Let $U=X\setminus X_{s}$ and let $V=Y\setminus Y_{s}$. Since $Y_{s}$ is closed in $Y$ and $f(X_{s})\subseteq Y_{s}$, the closure of $f(X_{s})$ cannot contain the generic points of $V$. Since $f$ is dominant its restriction $f:U\rightarrow V$ is still dominant. Now both $U$ and $V$ are Jacobson by [8, Lemma 3.18 (1)] and [8, Lemma 3.14] implies that $\dim V\leq\dim U$. Let $Z$ be the closure of $U$ in $X$ and let $W$ be the closure of $V$ in $Y$. The assumption (2) and Lemma [8, Lemma 3.18 (5)] imply that $\dim Z=\dim U+1$. The proof of [8, Lemma 3.18 (5)], namely the inequality (14) in [8], implies that $\dim W\leq\dim V+1$. Thus $\dim W\leq\dim Z$.
@@
To prove the claim we observe that the map $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\rightarrow G^{0}\times X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$, $x\mapsto(1,x)$ induces an isomorphism $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\sslash H^{0}\overset{\cong}{\longrightarrow}(G^{0}\times^{H^{0}}X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}})\sslash G^{0}$ with the inverse map induced by $(g,x)\mapsto x$. Now $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\sslash H^{0}=X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}$ is a spectrum of a finite product of local noetherian $\mathcal{O}$-algebras by (31). Since $G^{0}$ is connected every irreducible component $X^{\prime}$ of $X$ is $G^{0}$-invariant by [8, Lemma 2.1] and hence its image in $X\sslash G^{0}$ is closed by Corollary 2 (ii) to [52, Proposition 9]. Thus the image of $X^{\prime}$ in $X\sslash G^{0}=X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}$ will contain a closed point. Since $X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}\rightarrow X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is finite, it sends closed points to the closed point of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. This implies that $X$ satisfies the assumption (2) of Lemma 14.2. The assumption (1) holds as $Y$ is by definition the closure of $f(X)$. The argument in the special fibre is the same. ∎
@@
If $x\in\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ then $\rho_{x}(\Gamma_{F})$ is contained in $P(\kappa(x))$ for some R-parabolic subgroup $P$ of $G_{\kappa(x)}$ such that R-Levi of $P$ is conjugate to $L_{\kappa(x)}$. Thus if $x\not\in G^{0}\cdot\overline{X}^{\mathrm{gen}}_{L,\overline{\rho}^{\mathrm{ss}}}$ then $\rho_{x}(\Gamma_{F})$ is not contained in any R-Levi subgroup of $P$, and so part (1) of Proposition 11.5 holds for $\rho_{x}$. Moreover, Proposition 14.4 implies that
@@
We claim that $\overline{X}^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$ locus is contained in the union of $Y_{\overline{\rho}^{\mathrm{ss}}}$, $\overline{Z}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$, $\overline{Z}_{LG,\overline{\rho}^{\mathrm{ss}}}$, for $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ of $G$ containing a fixed maximal split torus $T$ and satisfying $\dim G-\dim L>2$, and $\overline{D}_{L}$ for $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ of $G$ containing $T$ and satisfying $\dim G-\dim L=2$. Since this is a finite union and each piece has codimension of at least $1+[F:\mathbb{Q}_{p}]$ by Corollary 12.18, Propositions 13.18, 13.17, 14.5 we obtain the assertion. Moreover, if there are no $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ with $\dim G-\dim L=2$ then the same references imply that each piece has codimension at least $2[F:\mathbb{Q}_{p}]$.
We will now prove the claim. Let $x\in\overline{X}^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$ be a geometric point. If $x$ is not contained in $Y_{\overline{\rho}^{\mathrm{ss}}}$ then it is contained in $\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ for some $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi $L$ of $G$ containing $T$. If $L=G$ then $x\in\overline{Z}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ by definition of the irreducible special locus. If $L\neq G$ then $\dim G-\dim L\geq 2$ by part (3) of Lemma 9.11. If $\dim G-\dim L>2$ then $x\in\overline{Z}_{LG,\overline{\rho}^{\mathrm{ss}}}$, as this locus contains $\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$. If $\dim G-\dim L=2$ then Proposition 14.5 implies that $x\in\overline{D}_{L}$. ∎

### v2

@@ 13.4. Consequences
Let $x$ be either a closed point of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}\setminus Y_{\overline{\rho}^{\mathrm{ss}}}$ or a closed point of $Y_{\overline{\rho}^{\mathrm{ss}}}$ and let $\Lambda$ be a coefficient ring for $\kappa(x)$ introduced in Section 3. Then the following hold:
@@
    $R^{\square}_{G,\rho_{x}}$ is a flat $\Lambda$-algebra of relative dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$ and is a complete intersection;
@@
    if $\mathrm{char}(\kappa(x))=p$ then $R^{\square}_{G,\rho_{x}}/\varpi$ is a complete intersection of dimension $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$.
@@
The proof is essentially the same as the proof of [6, Corollary 3.38], so we only give a sketch. Corollary 3.8 and Lemma 3.9 applied with $A=\Lambda$ and $B=R^{\square}_{G,\rho_{x}}$ imply that to prove part (1) it is enough to show that
@@
If $\mathrm{char}(\kappa(x))=0$ then $\Lambda=\kappa(x)$. If $\mathrm{char}(\kappa(x))=p$ then $\Lambda$ is a DVR with uniformiser $\varpi$ and the fibre ring in (80) is $R^{\square}_{G,\rho_{x}}/\varpi$. Thus, if we can prove (80) then Lemma 3.9 applied with $A=\kappa(x)$ and $B=R^{\square}_{\rho_{x}}/\varpi$ will imply part (2).
@@
The proof is essentially the same as the proof of [6, Corollary 3.40] so we only give a sketch. Lemma 5.17 relates $R^{\square}_{G,\rho_{x}}$ and the local ring $\mathcal{O}_{X,x}$, where $X:=X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ and $x$ is a closed point of $X$. Using Corollary 13.27 we deduce that $\mathcal{O}_{X,x}$ are complete intersection and $\mathcal{O}$-flat for all closed points $x$. Thus $X$ is flat over $\operatorname{Spec}\mathcal{O}$ and locally complete intersection. The last property implies that $X$ is equidimensional. Part (1) implies part (2). ∎
@@
We will prove (1); the proof of (2) is identical. Let $\eta$ be a generic point of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Let $\bar{\eta}$ be a geometric point above $\eta$. It follows from [57, Theorem 3 (ii)] that there is a geometric point $\xi:\operatorname{Spec}\kappa(\bar{\eta})\rightarrow X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ above $\bar{\eta}$. Let $X$ be an irreducible component of $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ containing $\xi$ and let $X^{\mathrm{git}}$ be the image of $X$ in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Then $X^{\mathrm{git}}$ contains $\eta$. It follows from [6, Lemma 2.1] that $X$ is $G^{0}$-invariant. Since $X$ is closed in $X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$, [57, Theorem 3 (iii)] implies that $X^{\mathrm{git}}$ is closed in $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. Since $X$ is irreducible, $X^{\mathrm{git}}$ is also irreducible, [59, Tag 0379]. We conclude that $X^{\mathrm{git}}$ is the irreducible component of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ containing $\eta$.
@@
Corollaries 13.31 and 13.32 are the analogues of the main results of [7]; compare [7, Theorem 5.5.1]. Even if $G=\operatorname{GL}_{d}$ the proof given here is new: although we use many techniques introduced by Böckle and Juschka in [7], we never appeal to [7, Section 5].
@@
Using Lemma 3.10 it is enough to prove the statement after replacing $\kappa$ by a finite extension. Using Proposition 6.6 we may assume that there exist $g\in G^{0}(\kappa)$ such that $\rho^{\prime}(\gamma):=g\rho(\gamma)g^{-1}\in G(\mathcal{O}_{\kappa})$ for all $\gamma\in\Gamma_{F}$. Since $G$ is smooth over $\mathcal{O}$ we may pick $h\in G(\Lambda)$ which maps to $g$ in $G(\kappa)$. Conjugation by $h$ induces an isomorphism of $\Lambda$-algebras between $R^{\square}_{G,\rho}$ and $R^{\square}_{G,\rho^{\prime}}$, so we may assume that $\rho$ takes values in $G(\mathcal{O}_{\kappa})$. Let $\overline{\rho}$ be the reduction of $\rho$ modulo the uniformiser of $\mathcal{O}_{\kappa}$. Let $\mathcal{O}$ be the Witt ring of residue field of $\mathcal{O}_{\kappa}$. Since $\rho$ is a deformation of $\overline{\rho}$ to $\mathcal{O}_{\kappa}$ we obtain a map of local $\mathcal{O}$-algebras $R^{\square}_{G,\overline{\rho}}\rightarrow\mathcal{O}_{\kappa}$. Lemma 3.11 says that $R^{\square}_{G,\overline{\rho}}$ is naturally isomorphic to the completion of $\Lambda\otimes_{\mathcal{O}}R^{\square}_{G,\overline{\rho}}$ with respect to the kernel $\mathfrak{q}$ of the natural map to $\kappa(x)$. Corollary 13.27 together with [6, Lemmas 3.36, 3.37] allow us to bound the dimension of the completion by $\dim\Lambda+\dim G_{k}([F:\mathbb{Q}_{p}]+1)$, and if $\mathrm{char}(\kappa(x))=p$ then the dimension of the special fibre of the completion by $\dim G_{k}([F:\mathbb{Q}_{p}]+1)$. We obtain $\dim\kappa\otimes_{\Lambda}R^{\square}_{G,\rho}\leq\dim G_{k}([F:\mathbb{Q}_{p}]+1)$ and the proof of Corollary 13.27 goes through. ∎
@@
Every continuous representation $\overline{\rho}:\Gamma_{F}\rightarrow G(k)$ can be lifted to characteristic zero. More precisely, there is a finite extension $L^{\prime}$ of $L$ with a ring of integers $\mathcal{O}^{\prime}$ and uniformiser $\varpi^{\prime}$ and a continuous representation $\rho:\Gamma_{F}\rightarrow G(\mathcal{O}^{\prime})$ such that $\rho\equiv\overline{\rho}\pmod{\varpi^{\prime}}$.
@@
The proof is essentially the same as the proof of [6, Proposition 3.55]. The map $X^{\square}_{G,\overline{\rho}}\rightarrow X^{\mathrm{gen}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is flat, since it is a localisation followed by completion. Let $V$ be a preimage of $V^{\operatorname{n-spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ in $X^{\square}_{G,\overline{\rho}}$. Then $V$ is contained in $X^{\square}_{G,\overline{\rho}}\setminus\{\ast\}$ and Lemma 3.54 in [6] together with Corollary 13.30 implies that $V$ is Zariski dense in $X^{\square}_{G,\overline{\rho}}$. The set of closed points in $V$ is equal to $\Sigma^{\operatorname{n-spcl}}$. Since $V$ is open $X^{\square}_{G,\overline{\rho}}\setminus\{\ast\}$, it is Jacobson, and hence $\Sigma^{\operatorname{n-spcl}}$ is dense in $V$ and also in $X^{\square}_{G,\overline{\rho}}$. The same argument works for the special fibre. ∎
@@ 14. Non-special locus
Let $H$ be a closed generalised reductive subgroup of $G$ such that $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ is non-empty. Then $H^{0}$ acts on $G^{0}\times X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ via $h\cdot(g,x)=(gh^{-1},h\cdot x)$. We let $G^{0}\times^{H^{0}}X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$ be the GIT quotient for this action. Since the action is free the fibre at a geometric point $x$ is isomorphic to $H^{0}_{\kappa(x)}$. This implies that
@@
The following Lemma formalises an argument made in [6, Proposition 3.8].
@@
Let $U=X\setminus X_{s}$ and let $V=Y\setminus Y_{s}$. Since $Y_{s}$ is closed in $Y$ and $f(X_{s})\subseteq Y_{s}$, the closure of $f(X_{s})$ cannot contain the generic points of $V$. Since $f$ is dominant its restriction $f:U\rightarrow V$ is still dominant. Now both $U$ and $V$ are Jacobson by [6, Lemma 3.18 (1)] and [6, Lemma 3.14] implies that $\dim V\leq\dim U$. Let $Z$ be the closure of $U$ in $X$ and let $W$ be the closure of $V$ in $Y$. The assumption (2) and [6, Lemma 3.18 (5)] imply that $\dim Z=\dim U+1$. The proof of [6, Lemma 3.18 (5)], namely the inequality (14) in [6], implies that $\dim W\leq\dim V+1$. Thus $\dim W\leq\dim Z$.
@@
To prove the claim we observe that the map $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\rightarrow G^{0}\times X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}$, $x\mapsto(1,x)$ induces an isomorphism $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\sslash H^{0}\overset{\cong}{\longrightarrow}(G^{0}\times^{H^{0}}X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}})\sslash G^{0}$ with the inverse map induced by $(g,x)\mapsto x$. Now $X^{\mathrm{gen}}_{H,\overline{\rho}^{\mathrm{ss}}}\sslash H^{0}=X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}$ is a spectrum of a finite product of local noetherian $\mathcal{O}$-algebras by (31). Since $G^{0}$ is connected every irreducible component $X^{\prime}$ of $X$ is $G^{0}$-invariant by [6, Lemma 2.1] and hence its image in $X\sslash G^{0}$ is closed by Corollary 2 (ii) to [57, Proposition 9]. Thus the image of $X^{\prime}$ in $X\sslash G^{0}=X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}$ will contain a closed point. Since $X^{\mathrm{git}}_{H,\overline{\rho}^{\mathrm{ss}}}\rightarrow X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is finite, it sends closed points to the closed point of $X^{\mathrm{git}}_{G,\overline{\rho}^{\mathrm{ss}}}$. This implies that $X$ satisfies the assumption (2) of Lemma 14.2. The assumption (1) holds as $Y$ is by definition the closure of $f(X)$. The argument in the special fibre is the same. ∎
@@
If $x\in\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ then $\rho_{x}(\Gamma_{F})$ is contained in $P(\kappa(x))$ for some R-parabolic subgroup $P$ of $G_{\kappa(x)}$ such that some R-Levi of $P$ is conjugate to $L_{\kappa(x)}$. Thus if $x\not\in G^{0}\cdot\overline{X}^{\mathrm{gen}}_{L,\overline{\rho}^{\mathrm{ss}}}$ then $\rho_{x}(\Gamma_{F})$ is not contained in any R-Levi subgroup of $P$, and so part (1) of Proposition 11.5 holds for $\rho_{x}$. Moreover, Proposition 14.4 implies that
@@
We claim that $\overline{X}^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$ is contained in the union of $Y_{\overline{\rho}^{\mathrm{ss}}}$, $\overline{Z}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$, $\overline{Z}_{LG,\overline{\rho}^{\mathrm{ss}}}$, for $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ of $G$ containing a fixed maximal split torus $T$ and satisfying $\dim G-\dim L>2$, and $\overline{D}_{L}$ for $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ of $G$ containing $T$ and satisfying $\dim G-\dim L=2$. Since this is a finite union and each piece has codimension of at least $1+[F:\mathbb{Q}_{p}]$ by Corollary 12.18, Propositions 13.18, 13.17, 14.5 we obtain the assertion. Moreover, if there are no $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi subgroups $L$ with $\dim G-\dim L=2$ then the same references imply that each piece has codimension at least $2[F:\mathbb{Q}_{p}]$.
We will now prove the claim. Let $x\in\overline{X}^{\operatorname{spcl}}_{G,\overline{\rho}^{\mathrm{ss}}}$ be a geometric point. If $x$ is not contained in $Y_{\overline{\rho}^{\mathrm{ss}}}$ then it is contained in $\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$ for some $\overline{\rho}^{\mathrm{ss}}$-compatible R-Levi $L$ of $G$ containing $T$. If $L=G$ then $x\in\overline{Z}^{\operatorname{spcl}}_{GG,\overline{\rho}^{\mathrm{ss}}}$ by definition of the absolutely irreducible special locus. If $L\neq G$ then $\dim G-\dim L\geq 2$ by part (3) of Lemma 9.11. If $\dim G-\dim L>2$ then $x\in\overline{Z}_{LG,\overline{\rho}^{\mathrm{ss}}}$, as this locus contains $\overline{{V}}_{LG,\overline{\rho}^{\mathrm{ss}}}$. If $\dim G-\dim L=2$ then Proposition 14.5 implies that $x\in\overline{D}_{L}$. ∎